# Actions Speak Louder Than Words: Rate-Reward Trade-off in Markov Decision Processes

### v1

@@ 1 Introduction
Motivated by these challenges, this paper explores implicit communication in a Markov Decision Process (MDP) environment—communication through actions—which facilitates information transmission from the MDP controller to other agents that can observe the MDP states. Effectively utilizing this internal channel has the potential to reduce the dependence on dedicated communication channels. However, these communication capabilities come with inherent trade-offs. As we will demonstrate, using this channel for communication often leads to a degradation in MDP control performance. This raises a fundamental challenge in balancing control and communication objectives, underscoring the need for a cohesive design that integrates both control and communication.
@@
An MDP consists of a controller and an environment (Puterman, 2014). At each time step, the environment is in some state $s$, and the controller selects some action $a$. Upon executing action $a$, the environment stochastically transitions to a new state $s^{\prime}$ and generates a reward. The controller’s objective is to find an optimal policy for selecting an action at each time to maximize its accumulated reward over a given time horizon. Consider that the controller now wants to communicate with another agent in the same environment that can also observe the environment state.
@@
In communication theory, a finite-state channel (FSC) (Gallager, 1968) is an input-output system with states, where the output depends on both the input and the current state. Messages are encoded into the input sequence, resulting in a corresponding output sequence. The receiver decodes the message from this output sequence. From this perspective, the state transition of an MDP from $s$ to $s^{\prime}$ upon taking action $a$ can be viewed as an FSC from the controller to the agent. We refer to this internal channel within the MDP as an action-state channel. To communicate through the action-state channel, we need an encoder that maps the message to a sequence of actions, and a decoder that translates the resulting state sequence back into the original message. However, the objective of this encoder differs from that of the controller: the encoder aims to maximize the transmission rate and reliability of its message, while the controller seeks to maximize the accumulated reward. These two objectives are generally inconsistent, and a trade-off between the two must be sought.
@@
In this paper, we first investigate the trade-off between the capacity (i.e., the maximum achievable transmission rate) of the action-state channel and the MDP reward in the infinite horizon regime. We demonstrate that the capacity of this channel can be expressed in a simple form—as the conditional mutual information between the input and output conditioned on the channel state. We also show that the capacity-reward trade-off can be characterized by a convex optimization, which can be solved numerically. While the capacity-reward trade-off provides an upper bound on the practically achievable rate under certain reward constraints, solving it does not yield a practical coding scheme. We then propose a practical framework for the integrated control and communication task in the finite block-length regime. The challenge of designing such a framework is twofold: (1) balancing the control and communication performance, and (2) dealing with the non-differentiability of the action-state channel. To tackle these issues, we propose Act2Comm, a transformer-based coding scheme in which encoder and decoder are trained iteratively.
@@ 2 Related work
Sokota et al. (2022) explored a similar concept of communication via MDPs. In their study, the receiver can observe the entire trajectory, including both the action and state sequences. This enables the controller to encode (compress) messages into the action sequence, and the receiver can subsequently decode the messages from the trajectory. Essentially, this is a randomized source coding problem. However, in most practical scenarios, while the MDP state is a physical signal observable by the receiver, the controller’s actions are usually not observable to other agents. Therefore, in our work, we assume the receiver can only observe the state sequence. This shifts the problem from source coding to channel coding. Karabag et al. (2019) also examined a similar system, but their focus was on developing policies that restrict the observer’s ability to infer transition probabilities.
FSC represents a general class of communication channels, and its study has been a long-standing problem in information and coding theory. Blackwell et al. (1958) studied the capacity of indecomposable FSCs without feedback. Subsequent studies in the non-feedback setting include Verdu & Han (1994) and Goldsmith & Varaiya (1996). The capacity of FSCs with feedback was examined by Massey (1990) and Permuter et al. (2009). More recently, Shemuel et al. (2022) explored the capacity of FSCs with feedback and state information at the encoder. However, these results express capacity in multi-letter forms, relying on the entire input and output sequences as their lengths approach infinity. Although Sabag et al. (2017) provided a single-letter upper bound for the feedback capacity of unifilar FSCs, exact single-letter expressions for FSC capacity are generally unknown. The action-state channel studied in this paper is a special FSC with state and feedback at the encoder. Utilizing the unique structure of this channel, we derive a single-letter expression for its capacity.
@@
Notations: For any $x_{t}$ with $t\geq 1$, $\bm{x_{i}^{k}}$ denotes the sequence $\{x_{i},x_{i+1},\ldots,x_{k}\}$, where $\bm{x_{1}^{k}}$ is written as $\bm{x^{k}}$. $|\mathcal{X}|$ denotes the cardinality of the set $\mathcal{X}$. A detailed notation table is provided as Table. 1.
@@ 3 Preliminaries and System Model
An MDP can be characterized by a tuple $(\mathcal{S},\mathcal{X},\bm{T},r,\alpha)$, where $\mathcal{S}$ is the state space, $\mathcal{X}$ is the action space, $\bm{T}$ is the transition kernel, $r:\mathcal{S}\times\mathcal{X}\rightarrow\mathbb{R}$ is the bounded reward function, and $\alpha$ is the initial state distribution. At each time step $t$, taking action $x_{t}$ in state $s_{t}$ results in a reward $r(x_{t},s_{t})$ and a state transition from $s_{t}$ to $s_{t+1}$, where $s_{t+1}$ is sampled from the distribution $\bm{T}(\cdot|s_{t},x_{t})$. We assume both $\mathcal{S}$ and $\mathcal{X}$ are finite sets. A stationary deterministic policy is a mapping $\pi:\mathcal{S}\to\mathcal{X}$ that selects action $x_{t}$ based on state $s_{t}$ at each time $t$. The objective is to find an optimal policy that maximizes the long-term average reward, as follows:
@@
To transmit a message $m$, an encoder generates a sequence of channel inputs $\bm{x}^{t}$ as codewords. When each $x_{t}$ is input to the channel, the channel transitions to a new state $s_{t+1}$ and produces an output $y_{t}$. The decoder then collects the output sequence $\bm{y}^{t}$ to reconstruct $m$. We suppose that the channel state is available to the encoder but not to the decoder, and that the outputs are fed back to the encoder.
Let $\mathcal{M}$ denote the set of messages, with each message $m$ uniformly sampled from $\mathcal{M}$. The encoder is defined as a sequence of mappings, $\mathcal{E}\triangleq\{\mathcal{E}_{1},\ldots,\mathcal{E}_{n}\}$, where each mapping $\mathcal{E}_{t}:\mathcal{M}\times{\mathcal{S}}^{t}\times\mathcal{X}^{t-1}\times{\mathcal{{Y}}}^{t-1}\to{\mathcal{X}}$ generates the channel input at time $t$. In other words, the channel input at time $t$, $x_{t}=\mathcal{E}_{t}({m},\bm{s^{t}},\bm{x^{t-1}},\bm{{y}^{t-1}})$, is a function of $m$ and all the historical information available at the transmitter up to time $t$. The decoder is defined as the mapping, $\mathcal{D}:\mathcal{Y}^{n}\to{\mathcal{M}}$, which reconstructs the message from all $n$ channel outputs, $\hat{m}=\mathcal{D}(\bm{y}^{n})$. The pair $(\mathcal{E},\mathcal{D})$ constitutes a code, where $n$ represents the code length. Suppose the message set is =$\mathcal{M}=\{1,2,\ldots,2^{k}\}$, then each message can be represented with $k$ bits. The rate of the code $(\mathcal{E},\mathcal{D})$ is defined as $R_{\mathcal{E},\mathcal{D}}=k/n$.
The probability of error $P_{e}^{(n)}$ for $(\mathcal{E},\mathcal{D})$ is defined as $P_{e}^{(n)}=\Pr(\mathcal{D}(\bm{y}^{n})\neq m|m\text{ is sent})$. A rate $R$ is deemed achievable if there exists a code $(\mathcal{E},\mathcal{D})$ such that the error probability of the transmission approaches zero as $n\to\infty$. Consequently, the capacity of the FSC is defined as the supremum of all achievable rates. In other words, channel capacity reveals the maximum rate required for error-free transmission when the code length approaches infinity. In practice, however, constructing a code with infinite code length is unfeasible. Hence practical channel coding aims to balance the trade-off between $R$ and $P_{e}^{(n)}$ with a finite code length $n$. For instance, we design codes to maximize the rate while ensuring that the probability of error remains below a certain threshold $\sigma>0$:
@@
(b) Controller & Encoder: Within this framework, the MDP controller and the FSC encoder represent the two aspects of the same entity, jointly responsible for selecting an action $x_{t}\in\mathcal{X}$ at each time step. However, their objectives differ: the controller aims to maximize the reward, while the FSC encoder seeks to maximize the message rate. Unlike the controller, which can focus on stationary deterministic policies, the encoder must account for more complex policy forms. For any message $m$, the encoder described in the previous part can be viewed as a history-dependent policy for the MDP. Therefore, we consider the joint control and coding policy in its most general form. The policy $\mathcal{E}$ is represented as a sequence of mappings $\{\mathcal{E}_{i}:0\leq i\leq n-1\}$, where $\mathcal{E}_{i}$ is defined as $\mathcal{E}_{i}:\mathcal{M}\times{\mathcal{S}}^{i}\times\mathcal{X}^{i-1}\to{\mathcal{X}}$. Each message is transmitted via a sequence of $n$ actions. For example, if the controller begins to transmit a message $m$ at time $t$, then $x_{t+i}=\mathcal{E}_{i}(m,\bm{s}_{t}^{t+i},\bm{x}_{t}^{t+i-1})$ for $10\leq i\leq n-$ 1. We assume the controller always has a new message ready for transmission immediately after completing the transmission of the previous message, and each message is uniformly sampled from $\mathcal{M}$. The long-term average reward of the MDP under policy $\mathcal{E}$ is denoted by $G_{\mathcal{E}}$.
@@ 4 The Capacity-Reward Trade-off
In this section, we analyze the trade-off between the capacity of the action-state channel and the MDP reward. While the results may not offer direct guidance for practical coding—since the capacity is typically achievable only in the infinite-horizon regime (i.e., when the code length $n\to\infty$)—they delineate the fundamental performance limits of communication via actions in MDPs, thus holding substantial theoretical importance. All proofs are detailed in Appendix B.
@@
As previously discussed, a general encoder for an FSC generates a channel input based on all the historical state and feedback information. However, Theorem 1 reveals a surprising fact: the capacity of the action-state channel can be achieved by encoding messages into a stationary randomized policy for the MDP, without relying on historical information.
@@
Denote by $C(V)$ the capacity of the action-state channel with reward constraint $V$. We have:
@@
for any $w_{n}\in\mathcal{W}$, $s\in\mathcal{S}$, and $x\in\mathcal{X}$. The gradient $\partial I/\partial w=[\nabla I_{w}(s,x)]_{s,x}$ then can be used in the gradient descent method to solve the optimization problem in Theorem 2.
@@ 5 Act2Comm: A practical coding scheme
Let $\mathcal{U}\triangleq\mathcal{X}^{|\mathcal{S}|}$, where each ${u}\in\mathcal{U}$ is referred to as a decision rule because the $i$-th element of $u$ can be viewed as an action prescribed for state $i$. A deterministic control policy can be defined as a sequence $\{u_{t}:t\geq 1\}$, where $u_{t}\in\mathcal{U}$ is the decision rule at time $t$. To facilitate block coding, we first convert the action-state channel into an extended action-state (EAS) channel (Fig. 6 in Appendix) using Shannon’s method (Shannon, 1958). In particular, we conceptually separate the encoder and controller, considering the controller as an integral component of the EAS channel. The channel state is assumed to be available at the controller but not the encoder. At each time $t$, the encoder selects a decision rule ${u_{t}}$ from $\mathcal{U}$. Then the controller uses ${u_{t}}$ and the state $s_{t}$ to determine an action $x_{t}={u_{t}}(s_{t})$. Consequently, the EAS channel has an input alphabet $\mathcal{U}$, output alphabet $\mathcal{S}$, and channel law:
$P_{S^{+}|S,U}(s_{t+1}|s_{t},u_{t})=P_{S^{+}|S,X}(s_{t+1}|s_{t},u_{t}(s_{t}))=\bm{T}(s_{t+1}|s_{t},u_{t}(s_{t})).$ (11)
The EAS channel and the action-state channel are equivalent, and we will focus on the EAS channel to develop our coding scheme. This approach allows us to map a data block to a sequence of decision rules without knowing future states. The actions for subsequent time steps can then be determined using these decision rules when the states are revealed. In the rest of this section, we detail the Act2Comm framework in five main components.
@@
As depicted in Fig. 2, given a $k$-bit message $\bm{m}\in\{0,1\}^{k}$, the Act2Comm first encodes $\bm{m}$ into a belief map $\bm{Z}\in\mathbb{R}^{|\mathcal{S}|\times\frac{k}{R}}$. This $\bm{Z}$ is subsequently mapped into a codeword $\bm{U}\in\mathcal{X}^{|\mathcal{S}|\times\frac{k}{R}}$ by a quantizer. At each time step $t$, the controller selects an action $x_{t}=\bm{U}[s_{t},t]$ for state $s_{t}$, and transitions into a new state according to the channel law $P(s_{t+1}|\bm{x^{t}},\bm{s^{t}})$. After $\frac{k}{R}$ time steps, the receiver decodes the message based on the accumulated observations $\bm{s}\in\mathcal{S}^{\frac{k}{R}}$.
Figure 2: Left: Workflow diagram of the Act2Comm scheme, with the dashed line indicating the gradient flow. Right: Illustration of the iterative training strategy, incorporating a critic network.
@@
where $\mathcal{Q}:\mathbb{R}^{|\mathcal{S}|\times\frac{k}{R}}\rightarrow\mathcal{X}^{|\mathcal{S}|\times\frac{k}{R}}$ is the quantization operation that maps the coding result to the nearest action index in the action space $\mathcal{X}$, and each element of resultant codeword $x_{t}=U[s_{t},t]$ represents the selected action for state $s_{t}$ at time step $t$.
@@
To model the trade-off between control and communication, we utilize a weighted loss function: $\mathcal{L}_{all}=\mathcal{L}_{com}+\lambda\cdot\mathcal{L}_{cont}$. The communication loss $\mathcal{L}_{com}$ is defined as the cross-entropy between the predictions from a critic network and their corresponding ground-truth, which quantifies the message decoding accuracy. To ensure control performance, we aim to make the coding policy behave closely to the target policy. Therefore, $\mathcal{L}_{cont}$ measures the “distance” between the coding policy and the target policy.
Let $\pi$ denote the target policy, with $\pi(x|s)$ representing the probability of taking action $x$ in state $s$. Let $f_{U}(x|s)$ denote the frequency of selecting $x$ in state $s$ across all decision rules in $\bm{U}$. We then use the mean square error (MSE) between $\pi$ and $f_{U}$ to measure the control loss for its stability in experiments. However, $f_{U}$ is non-differentiable during the backpropagation as it is discrete. To address this issue, we estimate $f_{U}(x|s)$ using $\bm{Z}$ in equation 12. Let $\bm{e}$ denote the all-one row vector, and define $\Gamma_{Z}(\bm{T},s,x)\triangleq\text{Sigmoid}(\gamma\cdot(|\mathcal{X}|\text{Sigmoid}(\bm{Z}[s,:])-x\bm{e}))$. When $\gamma>0$ is sufficiently large, $\Gamma_{Z}(\bm{T},s,x)$ is a $(kR)$-dim vector with elements close to either 00 or $1$. Additionally, $\Gamma_{Z}(\bm{T},s,x)\bm{e}^{\top}$ approximates the number of elements in $\bm{U}[s,:]$ that are not less than $x$. We refer to $\gamma$ as the temperature parameter and estimate $f_{U}(x|s)$ for $x>0$ as follows:
@@
(a) Critic network As shown in Fig. 2, a critic network is introduced to estimate the gradient during gradient backpropagation for the encoder optimization, which views the EAS channel and decoder as an unknown environment. Before each update of the encoder, a critic network is trained over $s_{in}$ inner steps to predict the logits for the neighbor belief maps of a given $\bm{Z}$. For each inner step $k$, the network is trained to predict the corresponding logits $\bm{\hat{M_{k}}}$ as $\bm{\hat{M}_{ck}}$ based on the neighbor belief maps sampled from $\bm{Z_{k}}=\bm{Z}+\bm{W_{k}}$, where $\bm{W_{k}}\in\mathbb{R}^{|\mathcal{S}|\times\frac{k}{R}}\sim\mathcal{N}(0,\sigma_{w}^{2})$ is the Gaussian noise term for neighboring sampling during the $k$-th inner step. The MSE loss, denoted as $\mathcal{L}_{cn}$, is utilized between $\bm{\hat{M}_{ck}}$ and $\bm{\hat{M}_{k}}$ to train the critic network. With this design, we aim to obtain a precise critic network to estimate gradients from neighbors of $\bm{Z}$ in a given environment, thereby helping update the encoder. Note that this extra training cost is incurred only during the offline training process, this critic network will be removed during the inference phase, as detailed in Appendix C.3-C.4..
(b) Iterative updating strategy As outlined in Fig. 2, at each update step $i$, we first train a critic network $\bm{\phi_{s_{in}}^{i}}$ using $s_{in}$ inner steps to learn to estimate the gradient around the samples. Next, the encoder is updated to $\bm{\theta^{i+1}}$ using the frozen decoder parameters $\bm{\xi^{i}}$ and the learned gradient estimation. Subsequently, the decoder is directly optimized with the loss function to obtain new parameters $\bm{\xi^{i+1}}$, while keeping the encoder frozen at $\bm{\theta^{i+1}}$. This process iteratively alternates between encoder and decoder updates, freezing one while optimizing the other at each step.

### v2

@@ 1 Introduction
Motivated by these challenges, this paper explores implicit communication in a Markov Decision Process (MDP)—communication through actions—which facilitates information transmission from the MDP controller to other agents that can observe the MDP states. Effectively utilizing this internal channel has the potential to reduce the dependence on dedicated communication channels. However, such a communication capability comes with inherent trade-offs. As we will demonstrate, using this channel for communication often leads to a degradation in MDP control performance. This raises a fundamental challenge in balancing control and communication objectives, underscoring the need for a cohesive design that integrates control and communication.
@@
An MDP consists of a controller and an environment (Puterman, 2014). At each time step, the environment is in some state $s$, and the controller selects an action $a$. Upon executing action $a$, the environment stochastically transitions to a new state $s^{\prime}$ and generates a reward. Controller’s objective is to find an optimal policy for selecting an action at each time to maximize its accumulated reward over a given time horizon. Consider that the controller wants to communicate with another agent (i.e., the receiver) that can also observe the environment state.
@@
In communication theory, a finite-state channel (FSC) (Gallager, 1968) is an input-output system with states, where the output depends on both the input and the current state. Messages are encoded into the input sequence, resulting in a corresponding output sequence. The receiver decodes the message from this output sequence. From this perspective, the state transition of an MDP from $s$ to $s^{\prime}$ upon taking action $a$ can be viewed as an FSC from the controller to the receiver. We refer to this internal channel within the MDP as an action-state channel. To communicate through the action-state channel, we need an encoder that maps the message to a sequence of actions, and a decoder that translates the resulting state sequence back into the original message. However, the objective of this encoder differs from that of the controller: the encoder aims to maximize the transmission rate and reliability of its message, while the controller seeks to maximize the accumulated reward. These two objectives are generally inconsistent, and a trade-off between the two must be sought.
@@
In this paper, we first investigate the trade-off between the capacity (i.e., the maximum achievable transmission rate) of the action-state channel and the MDP reward in the infinite horizon regime. We demonstrate that the capacity of this channel can be expressed in a simple form—as the conditional mutual information between the input and output conditioned on the channel state. We also show that the capacity-reward trade-off can be cast as a convex optimization problem, which can be solved numerically. While the capacity-reward trade-off provides an upper bound on the practically achievable rate under certain reward constraints, solving it does not yield a practical coding scheme. We then propose a practical framework for the integrated control and communication task in the finite block-length regime. The challenge in designing such a framework is twofold: (1) balancing the control and communication performances; and (2) dealing with the non-differentiability of the action-state channel. To tackle these issues, we propose Act2Comm, a transformer-based coding scheme in which encoder and decoder are trained iteratively.
@@ 2 Related work
Sokota et al. (2022) explored a similar concept of communication via MDPs. In their study, the receiver can observe the entire trajectory, including both the action and state sequences. This enables the controller to encode (compress) messages into the action sequence, and the receiver can subsequently decode the messages from the trajectory. Essentially, this is a source coding problem. However, in most practical scenarios, while the MDP state is a physical signal observable by the receiver, the controller’s actions are typically not directly observable by other agents. Therefore, in our work, we assume the receiver can only observe the state sequence. This shifts the problem from source coding to channel coding. Karabag et al. (2019) also examined a similar system, but their focus was on developing policies that restrict the observer’s ability to infer transition probabilities.
FSC represents a general class of communication channels, and its study has been a long-standing problem in information and coding theory. Blackwell et al. (1958) studied the capacity of indecomposable FSCs without feedback. Subsequent studies in the non-feedback setting include (Verdu & Han, 1994) and (Goldsmith & Varaiya, 1996). The capacity of FSCs with feedback was examined by Massey (1990) and Permuter et al. (2009). More recently, Shemuel et al. (2022) explored the capacity of FSCs with feedback and state information at the encoder. However, these results express capacity in multi-letter forms, relying on the entire input and output sequences as their lengths approach infinity. Although Sabag et al. (2017) provided a single-letter upper bound for the feedback capacity of unifilar FSCs, exact single-letter expressions for FSC capacity are generally unknown. The action-state channel studied in this paper is a special FSC with state and feedback at the encoder. Utilizing the unique structure of this channel, we derive a single-letter expression for its capacity.
@@
Notations: For any sequence $\{x_{t}:t\geq 1\}$, $\bm{x_{i}^{k}}$ denotes the sub-sequence $\{x_{i},x_{i+1},\ldots,x_{k}\}$, where $\bm{x_{1}^{k}}$ is written as $\bm{x^{k}}$. $|\mathcal{X}|$ denotes the cardinality of the set $\mathcal{X}$. A detailed notation table is provided in Table 1.
@@ 3 Preliminaries and System Model
An MDP can be characterized by a tuple $(\mathcal{S},\mathcal{X},\bm{T},r,\alpha)$, where $\mathcal{S}$ is the state space, $\mathcal{X}$ is the action space, $\bm{T}$ is the transition kernel, $r:\mathcal{S}\times\mathcal{X}\rightarrow\mathbb{R}$ is the bounded reward function, and $\alpha$ is the initial state distribution. At each time step $t$, taking action $x_{t}$ in state $s_{t}$ results in a reward $r(s_{t},x_{t})$ and a state transition from $s_{t}$ to $s_{t+1}$, where $s_{t+1}$ is sampled from the distribution $\bm{T}(\cdot|s_{t},x_{t})$. We assume both $\mathcal{S}$ and $\mathcal{X}$ are finite sets. A stationary deterministic policy is a mapping $\pi:\mathcal{S}\to\mathcal{X}$ that selects action $x_{t}$ based on state $s_{t}$ at each time $t$. The objective is to find an optimal policy that maximizes the long-term average reward, as follows:
@@
To transmit a message $m$, an encoder generates a sequence of inputs $\bm{x}^{t}$ as a codeword. When each $x_{t}$ is input to the channel, the channel transitions to a new state $s_{t+1}$ and produces an output $y_{t}$. The decoder then collects the output sequence $\bm{y}^{t}$ to reconstruct $m$. We suppose that the channel state is available to the encoder but not to the decoder, and that the outputs are fed back to the encoder.
Let $\mathcal{M}$ denote the set of messages, with each message $m$ uniformly sampled from $\mathcal{M}$. The encoder is defined as a sequence of mappings, $\mathcal{E}\triangleq\{\mathcal{E}_{1},\ldots,\mathcal{E}_{n}\}$, where each mapping $\mathcal{E}_{t}:\mathcal{M}\times{\mathcal{S}}^{t}\times\mathcal{X}^{t-1}\times{\mathcal{{Y}}}^{t-1}\to{\mathcal{X}}$ generates the channel input at time $t$. In other words, the channel input at time $t$, $x_{t}=\mathcal{E}_{t}({m},\bm{s^{t}},\bm{x^{t-1}},\bm{{y}^{t-1}})$, is a function of $m$ and all the historical information available at the transmitter up to time $t$. The decoder is defined as the mapping, $\mathcal{D}:\mathcal{Y}^{n}\to{\mathcal{M}}$, which reconstructs the message from all $n$ channel outputs, $\hat{m}=\mathcal{D}(\bm{y}^{n})$. The pair $(\mathcal{E},\mathcal{D})$ constitutes a code, where $n$ is called the code length. Suppose the message set is =$\mathcal{M}=\{1,2,3,\ldots,2^{k}\}$, then each message can be represented with $k$ bits. The rate of the code $(\mathcal{E},\mathcal{D})$ is defined as $R_{\mathcal{E},\mathcal{D}}=k/n$.
The error probability $P_{e}^{(n)}$ for $(\mathcal{E},\mathcal{D})$ is defined as $P_{e}^{(n)}=\Pr(\mathcal{D}(\bm{y}^{n})\neq m|m\text{ is sent})$. A rate $R$ is deemed achievable if there exists a code $(\mathcal{E},\mathcal{D})$ such that the error probability of the transmission approaches zero as $n\to\infty$. Consequently, the capacity of the FSC is defined as the supremum of all achievable rates. In other words, channel capacity reveals the maximum rate available for error-free transmission when the code length approaches infinity. In practice, however, constructing a code with infinite code length is unfeasible. Hence practical channel coding aims to balance the trade-off between $R$ and $P_{e}^{(n)}$ with a finite code length $n$. For instance, we design codes to maximize the rate while ensuring that the probability of error remains below a certain threshold $\sigma>0$:
@@
(b) Controller & Encoder: Within this framework, the MDP controller and the FSC encoder represent the two aspects of the same entity, jointly responsible for selecting an action $x_{t}\in\mathcal{X}$ at each time step. However, their objectives differ: the controller aims to maximize the reward, while the encoder seeks to maximize the message rate. Unlike the controller, which can focus on stationary deterministic policies, the encoder must account for more complex policy forms. For any message $m$, the encoder described in the previous part can be viewed as a history-dependent policy for the MDP. Therefore, we consider the joint control and coding policy in its most general form. The policy $\mathcal{E}$ is represented as a sequence of mappings $\{\mathcal{E}_{i}:0\leq i\leq n-1\}$, where $\mathcal{E}_{i}$ is defined as $\mathcal{E}_{i}:\mathcal{M}\times{\mathcal{S}}^{i}\times\mathcal{X}^{i-1}\to{\mathcal{X}}$. Each message is transmitted via a sequence of $n$ actions. For example, if the controller begins to transmit a message $m$ at time $t$, then $x_{t+i}=\mathcal{E}_{i}(m,\bm{s}_{t}^{t+i},\bm{x}_{t}^{t+i-1})$ for $10\leq i\leq n-$ 1. We assume the controller always has a new message ready for transmission immediately after completing the transmission of the previous message, and each message is uniformly sampled from $\mathcal{M}$. The long-term average reward of the MDP under policy $\mathcal{E}$ is denoted by $G_{\mathcal{E}}$.
@@ 4 The Capacity-Reward Trade-off
In this section, we analyze the trade-off between the capacity of the action-state channel and the MDP reward. While the results may not offer direct guidance for practical coding—since the capacity is typically achievable only in the infinite-horizon regime (i.e., when the code length $n\to\infty$)—they delineate the fundamental performance limits of communication via actions in MDPs, thus holding substantial theoretical importance. All proofs of this section are detailed in Appendix B.
@@
As previously discussed, a general encoder for an FSC generates a channel input based on all the historical state and feedback information. However, Theorem 1 reveals a surprising fact: the capacity of the action-state channel can be achieved by encoding messages using a stationary randomized policy for the MDP, without relying on historical information.
@@
Denote by $C(V)$ the capacity of the action-state channel with reward constraint $V$.
@@
for any $w_{n}\in\mathcal{W}$, $s\in\mathcal{S}$, and $x\in\mathcal{X}$. The gradient $\partial I/\partial w=[\nabla I_{w}(s,x)]_{s,x}$ then can be used in the gradient ascent method to solve the optimization problem in Theorem 2.
@@ 5 Act2Comm: A practical coding scheme
Let $\mathcal{U}\triangleq\mathcal{X}^{|\mathcal{S}|}$. Each ${u}\in\mathcal{U}$ is referred to as a decision rule as the $i$-th element of $u$ can be viewed as an action prescribed for state $i$. A deterministic control policy can be defined as a sequence $\{u_{t}:t\geq 1\}$, where $u_{t}\in\mathcal{U}$ is the decision rule at time $t$. To facilitate block coding, we first convert the action-state channel into an extended action-state (EAS) channel (Fig. 6 in Appendix) using Shannon’s method (Shannon, 1958). We conceptually separate the encoder and controller, considering the controller as part of the EAS channel. Channel state is assumed to be available at the controller, but not the encoder. At each time $t$, the encoder selects a decision rule ${u_{t}}$ from $\mathcal{U}$. Then the controller uses ${u_{t}}$ and the state $s_{t}$ to determine an action $x_{t}={u_{t}}(s_{t})$. Consequently, the EAS channel has input alphabet $\mathcal{U}$, output alphabet $\mathcal{S}$, and channel law:
$P_{S^{+}|S,U}(s_{t+1}|s_{t},u_{t})=P_{S^{+}|S,X}(s_{t+1}|s_{t},u_{t}(s_{t}))\triangleq\bm{T}(s_{t+1}|s_{t},u_{t}(s_{t})).$ (11)
The EAS channel and the action-state channel are equivalent. We will focus on the former to develop our coding scheme. This approach allows us to map a data block to a sequence of decision rules without knowing the future states. Actions for subsequent time steps can then be determined using these decision rules when the states are revealed. In the rest of this section, we detail the Act2Comm framework, which consists of five components.
@@
As depicted in Fig. 2, given a $k$-bit message $\bm{m}\in\{0,1\}^{k}$, the Act2Comm first encodes $\bm{m}$ into a belief map $\bm{Z}\in\mathbb{R}^{|\mathcal{S}|\times\frac{k}{R}}$. This $\bm{Z}$ is subsequently mapped into a codeword $\bm{U}\in\mathcal{X}^{|\mathcal{S}|\times\frac{k}{R}}$ by a quantizer. At each time step $t$, the controller selects an action $x_{t}=\bm{U}[s_{t},t]$ for state $s_{t}$, and the channel transitions into a new state according to the channel law $P_{S^{+}|S,X}$. After $\frac{k}{R}$ time steps, the receiver decodes the message based on the accumulated observations $\bm{s}\in\mathcal{S}^{\frac{k}{R}}$.
Figure 2: (Left) Workflow diagram of the Act2Comm scheme, with the dashed line indicating the gradient flow. (Right) Illustration of the iterative training strategy, incorporating a critic network.
@@
where $\mathcal{Q}:\mathbb{R}^{|\mathcal{S}|\times\frac{k}{R}}\rightarrow\mathcal{X}^{|\mathcal{S}|\times\frac{k}{R}}$ is the quantization operation that maps the coding result to the nearest action index in the action space $\mathcal{X}$, and each element of resultant codeword, $x_{t}=\bm{U}[s_{t},t]$, represents the selected action for state $s_{t}$ at time step $t$.
@@
To model the trade-off between control and communication, we utilize a weighted loss function to train the encoder: $\mathcal{L}_{all}=\mathcal{L}_{com}+\lambda\mathcal{L}_{cont}$. The communication loss $\mathcal{L}_{com}$ is defined as the cross-entropy between the predictions from a critic network and their corresponding ground-truth, which quantifies the message decoding accuracy. To ensure control performance, we aim to make the coding policy behave closely to the target policy. Therefore, $\mathcal{L}_{cont}$ measures the “distance” between the coding policy and the target policy.
Let $\pi$ denote the target policy, with $\pi(x|s)$ representing the probability of taking action $x$ in state $s$. Let $f_{U}(x|s)$ denote the frequency of selecting $x$ in state $s$ across all decision rules in $\bm{U}$. We then use the mean square error (MSE) between $\pi$ and $f_{U}$ to measure the control loss for its stability in experiments. However, $f_{U}$ is non-differentiable during the backpropagation as it is discrete. To address this issue, we estimate $f_{U}(x|s)$ using $\bm{Z}$ in equation 12. Let $\bm{e}$ denote the all-one row vector, and define $\Gamma_{Z}(\bm{T},s,x)\triangleq\text{Sigmoid}(\gamma(|\mathcal{X}|\text{Sigmoid}(\bm{Z}[s,:])-x\bm{e}))$. When $\gamma>0$ is sufficiently large, $\Gamma_{Z}(\bm{T},s,x)$ is a $(kR)$-dim vector with elements close to either 00 or $1$. Additionally, $\Gamma_{Z}(\bm{T},s,x)\bm{e}^{\top}$ approximates the number of elements in $\bm{U}[s,:]$ that are not less than $x$. We refer to $\gamma$ as the temperature parameter and estimate $f_{U}(x|s)$ for $x>0$ as follows:
@@
(a) Critic network. As shown in Fig. 2, a critic network is introduced to estimate the gradient during gradient backpropagation for the encoder optimization, which views the EAS channel and decoder as an unknown environment. Before each update of the encoder, a critic network is trained over $s_{in}$ inner steps to predict the logits for the neighbor belief maps of a given $\bm{Z}$. For each inner step $k$, the network is trained to predict the corresponding logits $\bm{\hat{M_{k}}}$ as $\bm{\hat{M}_{ck}}$ based on the neighbor belief maps sampled from $\bm{Z_{k}}=\bm{Z}+\bm{W_{k}}$, where $\bm{W_{k}}\in\mathbb{R}^{|\mathcal{S}|\times\frac{k}{R}}\sim\mathcal{N}(0,\sigma_{w}^{2})$ is the Gaussian noise term for neighboring sampling during the $k$-th inner step. The MSE loss, denoted as $\mathcal{L}_{cn}$, is utilized between $\bm{\hat{M}_{ck}}$ and $\bm{\hat{M}_{k}}$ to train the critic network. With this design, we aim to obtain a precise critic network to estimate gradients from neighbors of $\bm{Z}$ in a given environment, thereby helping update the encoder. Note that this extra training cost is incurred only during the offline training process, this critic network will be removed during the inference phase, as detailed in Appendices C.3-C.4..
(b) Iterative updating strategy. As outlined in Fig. 2, at each update step $i$, we first train a critic network $\bm{\phi_{s_{in}}^{i}}$ using $s_{in}$ inner steps to learn to estimate the gradient around the samples. Next, the encoder is updated to $\bm{\theta^{i+1}}$ using the frozen decoder parameters $\bm{\xi^{i}}$ and the learned gradient estimation. Subsequently, the decoder is directly optimized with the loss function to obtain new parameters $\bm{\xi^{i+1}}$, while keeping the encoder frozen at $\bm{\theta^{i+1}}$. This process iteratively alternates between encoder and decoder updates, freezing one while optimizing the other at each step.